OpenStudy (anonymous):

I am having difficulty understanding a question on probability. A standard number cube is tossed. Find the probability. P(3 or odd)

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

ill help

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

weres the question?

OpenStudy (anonymous):

I posted it. But here I'll do it again. A standard number cube is tossed. Find the probability. P(3 or odd)

OpenStudy (alyssa_xo):

what's the p(3)? what's the p(odd)?

OpenStudy (anonymous):

P(3) = 1/6 P(odd) = 3/6 or 1/2

OpenStudy (alyssa_xo):

or in probability means addition

OpenStudy (anonymous):

Yeah. Or means to add

OpenStudy (alyssa_xo):

I'm just not sure if we want to include 3 twice considering it's odd as well

jimthompson5910 (jim_thompson5910):

here's what I would do first list out every possible outcome of the 6 sided number cube